OpenStudy (anonymous):

Factorize \(x^2+3xy+2y^2+4x+5y+3\) Into product of two linear factors.

OpenStudy (anonymous):

this is not a common way but u can make a quadratic in terms of x for example and solve for its roots

OpenStudy (anonymous):

Can you explain a little?

OpenStudy (anonymous):

ok when we have a quadratic like this :\[x^2+4x+3\]and we want to factor it...actually we're going to find its roots...after factoring we write it like this \[(x+1)(x+3)\]-1 and -3 are the roots of quadratic now we have\[ x^2+3xy+2y^2+4x+5y+3=0\]find the roots of this quadratic for x

OpenStudy (anonymous):

Wow! That looks complicated! But, 3xy, 2y^2, 5y ... What to do with them?

OpenStudy (anonymous):

I just edited the question. Please see if it is easier.

OpenStudy (anonymous):

\[x^2+(3y+4)x+2y^2+5y+3=0\]first step is to evaluate discriminant...

OpenStudy (anonymous):

its same

OpenStudy (anonymous):

it will be a pretty expression...i promise

OpenStudy (anonymous):

Δ = (3y+4)^2 - 4(2y^2 + 5y + 3) = y^2 + 4y + 4 = (y+2)^2 ?!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so what are the roots?

OpenStudy (anonymous):

\[x =\frac{ -(3y+4) \pm(y+2)}{2}\]\[ = -y -1 \ or \ -2y-3\]?

OpenStudy (anonymous):

done

OpenStudy (anonymous):

\[x^2+3xy+2y^2+4x+5y+3=?\]

OpenStudy (anonymous):

Wow! That's a faster method!

OpenStudy (anonymous):

(x+2y+3)(x+y+1)

OpenStudy (anonymous):

Here's the slow one :( x^2 + 3xy + 2y^2 = (x+2y)(x+y) x^2 + 3xy + 2y^2 + 4x + 5y + 3 = (x+2y+l)(x+y+m) = x^2 + 3xy + 2y^2 + (l+m)x + (l+2m)y + lm l+m = 4 l+2m = 5 lm = 3 => l = 3 ; m = 1 So, x^2 + 3xy + 2y^2 + 4x + 5y + 3 = (x+2y+3)(x+y+1) Thanks! I learnt a better method now!

OpenStudy (anonymous):

welcome :)